OpenStudy (anonymous):

g(x)=7x+10 g(x-2)+g(1) evaluate the function

OpenStudy (anonymous):

\[g(1)=7\times 1+10=17\] \[g(x-2)=7(x-2)+10\] \[g(x-2)+g(1)=7(x-2)+10+17\] then algebra